# Trajectory-Based Multi-Objective Hyperparameter Optimization for Model Retraining

### v1

@@ C.2 Problems and Benchmarks
LCBench [46]: This benchmark is designed to give insights on multi-fidelity optimization with learning curves for Auto Deep Learning. LCBench was originally developed upon tabular data. Benefitting from the surrogate implementation by HPOBench (under Apache License 2.0) [12] and YAHPO Gym (under Apache License 2.0) [32], we are allowed to observe the intermediate model performance for any feasible hyperparameter setting after each epoch. The maximum number of epochs for training MLP is set to $50$. We choose to minimize validation loss and training time, which are commonly considered in many MOHPO studies. For a demonstration, we focus on tuning the five hyperparameters (i.e., $d=5$) of MLP including learning rate in $[1\times 10^{-4},1\times 10^{-1}]$, momentum in $[0.10,0.99]$, weight decay in $[1\times 10^{-5},1\times 10^{-1}]$, maximum dropout rate in $[0.0,1.0]$, and maximum number of neurons in $[64,1024]$. LCBench utilizes diverse datasets from AutoML Benchmark [15] hosted on OpenML [41]. For experimentation, we select “blood-transfus”, “higgs”, “jungle-chess-2pcs”, “kc1”, and “banck-marketing”, each of which has at least two attributes and between 500 and 1,000,000 data points.
Appendix D Additional Numerical Results
This section presents additional numerical results comparing TMOBO with three enhanced multi-objective optimization methods. It is noteworthy that all the algorithms follow a similar routine in each iteration: upon sampling a new hyperparameter setting, the training procedure for the corresponding ML model starts to proceed by epochs. The difference is that TMOBO dynamically determines the stopping epoch during training while the others determine it before training begins. By tracking the number of times of model training, Figure 4 and 5 depicts the performance of each algorithm against the number of iterations on 20 synthetic problems analyzed in Section 6.1. The four subplots in each row represent synthetic problems derived from the same multi-objective benchmark but with different trajectory characteristics.
@@ Appendix D Additional Numerical Results
Figure 5: Average log Hypervolume difference against iterations for each algorithm on synthetic problems generated from DTLZ1, DTLZ2, and DTLZ7. Each algorithm runs for 20 independent trials. The shaded region indicates two standard errors of the mean.
Firstly, on variants of two relatively simple ZDT benchmarks (in Figure 4), TMOBO quickly converges to a small HV difference value after executing a small number of iterations, and it maintains the lowest average HV difference until the end. Among the alternatives, the HV-based methods, $q$NEHVI-T and $q$EHVI-T, perform better and can find solutions that are comparable to TMOBO on the first two instances of ZDT2; however, as the trajectory characteristics become more complex, both algorithms are dominated by TMOBO.
@@
On variants of DTLZ benchmarks (in Figure 5), the difference between TMOBO and alternatives becomes larger, especially on the rows of DTLZ1 and DTLZ7, while the alternatives show similar performances to each other. On the synthetic problems derived from DTLZ1, all the algorithms struggle to converge adequately within 150 iterations, primarily because the DTLZ1 benchmark has multiple local optima. However, through the trials, the HV difference values reached by TMOBO are significantly better than the alternatives. Additionally, the DTLZ7 benchmark has a disconnected Pareto-optimal front and it challenges the capability of optimizers to maintain a diverse set of solutions. Notably, TMOBO emerges as the best choice for handling this type of challenge, as the alternatives tend to stagnate at relatively high HV difference values at an early stage.
@@
Figure 6: Average log Hypervolume difference against iterations for each algorithm on five different hyperparameter tuning tasks. Each algorithm runs for 20 independent trials. The shaded region indicates two standard errors of the mean.
Finally, on the five LCBench hyperparameter tuning tasks analyzed in Section 6.2, we also show the performance of algorithms against iterations in Figure 6. Similar to the comparison results based on time, when given the same number of iterations, TMOBO consistently achieves the smallest HV difference values throughout. In other words, the Pareto-optimal fronts found by TMOBO are not only better converged but also more diversified compared to those found by the alternatives.
Appendix E Replication Strategy
@@
Replicated observations at a query $(\bm{x},t)$ might be needed to achieve adequate accuracy, especially in the presence of significant noises. While this particular scenario is not the focus of our current study, we outline a straightforward method to extend the proposed TMOBO to accommodate the requirement for replicated observations when the computational budget is sufficient. It is essential to recognize that with each noisy observation obtained at $(\bm{x},t)$, the training process must start from scratch and will also output noisy observations from $(\bm{x},1)$ up to $(\bm{x},t-1)$. As a consequence, careful maintenance of multiple training procedures for replications is needed to ensure algorithmic efficiency. The pseudo-code for TMOBO-$P$, a modified version of TMOBO, is presented in Algorithm 2.
Algorithm 2 Framework of TMOBO-$P$
@@
Figure 7 shows the performance of TMOBO-$P$ with $P=1$, $4$, $16$, and $64$ on the synthetic problems derived from ZDT1. TMOBO proposed in the main paper can be considered a special case of TMOBO-$P$ with $P=1$. This time, we add Gaussian noise to each objective with a standard deviation of $10\%$ of the objective range in order to emphasize the influence of noises. It can be observed that the performance of TMOBO-$P$ improves significantly in terms of the HV difference as more replications are allowed. This improvement is evident not only in rapid convergence during the initial stages but also in obtaining high-quality results at the end when comparing TMOBO with $P\geq 16$ to TMOBO with $P<16$. In the meantime, benefiting from a large number of replications, TMOBO-$64$ has stable performance and its standard error is relatively small.
Figure 7: Average log Hypervolume difference against iterations for each algorithm on synthetic problems generated from ZDT1 with $10\%$ noise level. Each algorithm runs for 20 independent trials. The shaded region indicates two standard errors of the mean.

### v2

@@ C.2 Problems and Benchmarks
LCBench [Zimmer et al., 2021]: This benchmark is designed to give insights on multi-fidelity optimization with learning curves for Auto Deep Learning. LCBench was originally developed upon tabular data. Benefitting from the surrogate implementation by HPOBench (under Apache License 2.0) [Eggensperger et al., 2021] and YAHPO Gym (under Apache License 2.0) [Pfisterer et al., 2022], we are allowed to observe the intermediate model performance for any feasible hyperparameter setting after each epoch. The maximum number of epochs for training MLP is set to $50$. We choose to minimize validation loss and training time, which are commonly considered in many MOHPO studies. For a demonstration, we focus on tuning the five hyperparameters (i.e., $d=5$) of MLP including learning rate in $[1\times 10^{-4},1\times 10^{-1}]$, momentum in $[0.10,0.99]$, weight decay in $[1\times 10^{-5},1\times 10^{-1}]$, maximum dropout rate in $[0.0,1.0]$, and maximum number of neurons in $[64,1024]$. LCBench utilizes diverse datasets from AutoML Benchmark [Gijsbers et al., 2019] hosted on OpenML [Vanschoren et al., 2014]. For experimentation, we select “blood-transfus”, “higgs”, “jungle-chess-2pcs”, “kc1”, and “banck-marketing”, each of which has at least two attributes and between 500 and 1,000,000 data points.
CNN: This benchmark provides insights into hyperparameter optimization for modern computer vision applications. The convolutional neural network model MobileNetV2 is implemented in PyTorch and optimized for both validation accuracy and training efficiency. Following standard computer vision practices, we optimize four key hyperparameters (i.e., $d=4$), including learning rate in $[1\times 10^{-4},1\times 10^{-1}]$, momentum in $[0.0,0.99]$, weight decay in $[1\times 10^{-5},1\times 10^{-2}]$, and batch size in $[128,512]$. The model is trained for up to 50 epochs on CIFAR-10 dataset, which consists of 60,000 32 $\times$ 32 color images across 10 classes. This configuration represents a more complex and computationally intensive optimization scenario compared to LCBench tasks, allowing us to assess our method’s scalability to challenging real-world deep learning applications.
Appendix D Sensitivity Analysis on Algorithm Configurations
In this section, we conduct a series of sensitivity analyses to demonstrate how TMOBO’s performance responds to its key algorithmic configurations. These configurations include the parameters for candidate search, Monte Carlo approximation, data augmentation strategy, and early stopping.
D.1 Parameters for Candidate Search
Regarding the candidate search, we examine both the number of candidates and the search radius. Intuitively, a larger candidate set increases the likelihood of getting promising hyperparameter settings; however, it can impose unnecessary computational costs at the same time. As shown in the third row of Figure D.2, using $100d$ to $200d$ candidates generally leads to a lower (better) overall distribution in the boxplots (where $d$ is the dimension of the domain of hyperparameter settings), especially when the objectives become more complex. Therefore, we choose a default setting of $100d$ candidates to balance effectiveness and efficiency. This choice is further supported by the statistical analysis in Table D.1, which indicates that our default setting is statistically comparable to or better than other settings, except in scenario Q-P where $200d$ outperforms.
For the search radius $r$, we can observe from the last row of Figure D.2 that a larger value of radius generally leads to a lower distribution in the boxplots; however, the marginal gains diminish gradually, and performance can even deteriorate beyond $r=0.2$ in some cases. This behaviour is expected because the radius controls the balance between exploration and exploitation, and a larger radius biases the algorithm toward exploration by increasing the chance of selecting hyperparameter settings further from the center, which can slow convergence. Statistical results in Table D.1 indicate that $r=0.2$ is significantly better than settings below 0.2 in half of the cases and comparable to higher settings across all cases. Consequently, we adopt it as our default.
D.2 Parameter for Monte Carlo Approximation
Regarding the Monte Carlo approximation, we showcase the algorithm’s performance with the number of MC samples $M$ being 32, 64, 128, 256, and 512. It can be observed from the second row of Figure D.2 that the algorithm’s performance is relatively insensitive to changes in $M$ within this range. While increasing the number of MC samples can improve the integration of predictive uncertainty, this improvement does not necessarily translate into better overall algorithm performance, but it raises the computational burden due to the expensive TEHVI acquisition function. Considering the added computational cost associated with significantly larger values of $M$, we believe that 128 samples provide a good balance, capturing sufficient information for accurate TEHVI approximation without incurring excessive computational expense. Moreover, the statistical tests in Table D.1 indicate that the algorithm performance using default setting of 128 is comparable to both lower and higher sample sizes across all tested cases.
D.3 Parameter for Data Augmentation Strategy
We first compare data augmentation strategies by evaluating the impact of retaining 5, 10, or 15 observations per trajectory in the GP model. This analysis helps us assess whether a careful selection of augmented observations can manage GP complexity effectively while making reasonable prediction for optimization. Our findings in Figure D.2 indicate that the algorithm’s performance is relatively insensitive to changes in this parameter across the four problems under study; in some cases, augmenting fewer data points even leads to a lower distribution in the boxplots. Considering the computational complexity of training a GP model, we maintain 10 observations per trajectory as our default setting, with the results in Table D.1 showing that this choice yields performance statistically comparable to other settings.
In parallel, we explore scalable GP models as an alternative to manual data augmentation. To this end, we implement a sparse GP with inducing points using the GPyTorch library [Gardner et al., 2018] and compare its performance against the data augmentation strategy. The Sparse GP (SGP) approach avoids the need for manual data curation and, as shown in Table D.1, typically achieves performance comparable to the 10-observation strategy in most cases. This indicates that both a careful selection of augmented observations and the use of sparse approximations are effective means of managing GP complexity in TMOBO. Moreover, as more algorithm iterations are executed, scalable GP models can serve as a suitable alternative.
Figure D.2: Sensitivity analysis on key parameters and components of TMOBO on synthetic problems generated from ZDT1 after 100 algorithm iterations. Each row varies one factor, namely (1) data augmentation strategy, using GP trained with 5, 10, or 15 observations per trajectory, or using Sparse GP (SGP); (2) number of MC samples $M\in\{32,64,128,256,512\}$; (3) number of candidates $q\in\{10d,50d,100d,150d,200d\}$, where $d$ is the problem dimension; and (4) search radius $r\in\{0.05,0.1,0.2,0.3,0.4\}$. Each box plot shows the logarithm of HV difference computed over 10 trials and is normalized to [0, 1]. (A lower value indicates better performance.)
Table D.1: Mean (standard deviation) of normalized log HV differences for key parameters and components of TMOBO on synthetic problems generated from ZDT1 after 100 algorithm iterations. For each parameter, performance is evaluated against the default algorithm setting (in bold) using Wilcoxon rank-sum test at a 95% confidence level. “$+$” sign indicates that the parameter option performs significantly better than the baseline, “--” indicates significantly worse performance, and $\approx$” denotes statistically comparable results.
ZDT1(M-M’) ZDT1(M-Q) ZDT1(M-P) ZDT1(Q-P)
Mean (std) Sign Mean (std) Sign Mean (std) Sign Mean (std) Sign
Data augmentation strategy 5 0.589 (0.277) $\approx$ 0.603 (0.172) $\approx$ 0.613 (0.244) $\approx$ 0.587 (0.190) $\approx$
10 0.528 (0.234) $\approx$ 0.659 (0.179) $\approx$ 0.602 (0.191) $\approx$ 0.628 (0.206) $\approx$
15 0.682 (0.219) n.a. 0.666 (0.240) n.a. 0.586 (0.214) n.a. 0.430 (0.230) n.a.
SGP 0.686 (0.282) $\approx$ 0.722 (0.203) $\approx$ 0.554 (0.111) $\approx$ 0.684 (0.157) --
Number of MC samples $M$ 32 0.493 (0.298) $\approx$ 0.597 (0.289) $\approx$ 0.496 (0.206) $\approx$ 0.593 (0.220) $\approx$
64 0.552 (0.147) $\approx$ 0.587 (0.182) $\approx$ 0.470 (0.153) $\approx$ 0.460 (0.197) $\approx$
128 0.417 (0.225) n.a. 0.583 (0.231) n.a. 0.486 (0.232) n.a. 0.603 (0.217) n.a.
256 0.734 (0.155) -- 0.616 (0.212) $\approx$ 0.542 (0.213) $\approx$ 0.498 (0.278) $\approx$
512 0.558 (0.154) $\approx$ 0.631 (0.136) $\approx$ 0.535 (0.183) $\approx$ 0.651 (0.241) $\approx$
[…]
Search radius $r$ 0.05 0.808 (0.162) -- 0.686 (0.181) -- 0.701 (0.225) -- 0.775 (0.152) --
0.1 0.529 (0.205) $\approx$ 0.536 (0.119) $\approx$ 0.593 (0.132) -- 0.484 (0.132) $\approx$
0.2 0.463 (0.120) n.a. 0.420 (0.167) n.a. 0.412 (0.190) n.a. 0.405 (0.175) n.a.
0.3 0.535 (0.097) $\approx$ 0.392 (0.162) $\approx$ 0.429 (0.135) $\approx$ 0.415 (0.129) $\approx$
0.4 0.458 (0.177) $\approx$ 0.532 (0.183) $\approx$ 0.363 (0.142) $\approx$ 0.311 (0.152) $\approx$
D.4 Parameter for Early Stopping
Regarding the early stopping mechanism, we present a separate sensitivity analysis in Figure D.3 and Table D.2, because the changes in $\beta$ affect the number of training epochs via early stopping, and it is more meaningful to compare the results after a fixed number of training epochs. As shown in Figure D.3, a moderate value of $\beta$ (between 0.1 and 0.3) generally results in a lower median in the boxplots; while $\beta=0.05$ or $0.4$ results in a higher median, especially in the first three cases. Intuitively, a lower $\beta$ accounts for less predictive uncertainty and may cause premature termination of training, thereby encouraging exploration of more hyperparameter settings; conversely, a higher $\beta$ is likely to prolong training along a given trajectory, leading to excessive exploitation and reduced exploration. This balance between exploitation and exploration is crucial for algorithm performance. The statistical results in Table D.2 show that algorithm performance using the default setting $\beta=0.2$ is statistically comparable to the other settings. Based on these insights, we recommend choosing a moderate $\beta$ value for a general TMOBO application.
Figure D.3: Sensitivity analysis on early stopping parameter $\beta\in\{0.5,1.0,2.0,3.0,4.0\}$ of TMOBO on synthetic problems generated from ZDT1 after 3000 training epochs. Each box plot shows the logarithm of HV difference computed over 10 trials and is normalized to [0, 1]. (A lower value indicates better performance.)
Table D.2: Mean (standard deviation) of normalized log HV differences for early stopping parameter $\beta$ of TMOBO on synthetic problems generated from ZDT1 after 3000 training epochs. For each parameter, performance is evaluated against the default algorithm setting (in bold) using Wilcoxon rank-sum test at a 95% confidence level. “$+$” sign indicates that the parameter option performs significantly better than the baseline, “--” indicates significantly worse performance, and $\approx$” denotes statistically comparable results.
ZDT1(M-M’) ZDT1(M-Q) ZDT1(M-P) ZDT1(Q-P)
Mean (std) Sign Mean (std) Sign Mean (std) Sign Mean (std) Sign
Early stopping ($\beta$) 0.5 0.802 (0.178) $\approx$ 0.658 (0.225) $\approx$ 0.606 (0.158) $\approx$ 0.498 (0.268) $\approx$
1.0 0.758 (0.156) $\approx$ 0.460 (0.238) $\approx$ 0.525 (0.138) $\approx$ 0.404 (0.229) $\approx$
2.0 0.735 (0.131) n.a. 0.496 (0.208) n.a. 0.502 (0.234) n.a. 0.530 (0.211) n.a.
3.0 0.751 (0.272) $\approx$ 0.459 (0.282) $\approx$ 0.525 (0.161) $\approx$ 0.481 (0.235) $\approx$
4.0 0.805 (0.115) $\approx$ 0.761 (0.198) -- 0.501 (0.159) $\approx$ 0.618 (0.210) $\approx$
Appendix E Numerical Experiments and Results
E.1 Additional Results for Section 6.1
In this subsection, we evaluate TMOBO and alternative algorithms on 20 synthetic problems derived from standard multi-objective benchmarks, as discussed in Section 6.1. Figures E.4 and E.5 illustrate the average log HV difference of each algorithm against the number of iterations. Each row of subplots represents problems generated from the same benchmark but with varying trajectory complexities.
On the ZDT benchmarks (Figure E.4), TMOBO achieves rapid convergence, reaching a low HV difference value within a few iterations. It maintains this advantage throughout the optimization process and outperforms the alternatives. While $q$NEHVI-T and $q$EHVI-T demonstrate competitive performance on the first two instances of ZDT2, their ability to handle increasing trajectory complexity is limited. In such cases, TMOBO significantly dominates these methods.
@@ Appendix E Numerical Experiments and Results
Figure E.4: Average log Hypervolume difference against iterations for each algorithm on synthetic problems generated from ZDT1 and ZDT2. Each algorithm runs for 20 independent trials. The shaded region indicates two standard errors of mean.
@@
In contrast, the variants of the DTLZ benchmarks (Figure E.5) present greater challenges. On synthetic problems derived from DTLZ1, all algorithms struggle to converge adequately within 150 iterations, primarily due to the existence of multiple local optima. However, TMOBO achieves noticeably better HV difference values, which highlights its robustness in handling complex landscapes. DTLZ7 benchmark challenges the capability of optimizers to maintain a diverse set of solutions as it has a disconnected Pareto-optimal front. Notably, TMOBO emerges as the best choice for handling this type of challenge, as the alternatives tend to stagnate at relatively high HV difference values at an early stage.
@@
Figure E.5: Average log Hypervolume difference against iterations for each algorithm on synthetic problems generated from DTLZ1, DTLZ2, and DTLZ7. Each algorithm runs for 20 trials. The shaded region indicates two standard errors of mean.
Figure E.6: Average log Hypervolume difference against iterations for each algorithm on five different hyperparameter tuning tasks. Each algorithm runs for 20 independent trials. The shaded region indicates two standard errors of the mean.
E.2 Additional Results for Section 6.2
Next, we assess TMOBO and alternative algorithms on five hyperparameter tuning tasks from LCBench, as described in Section 6.2. These tasks involve optimizing validation loss and training cost for an MLP model. Figure E.6 compares the algorithms in terms of average log HV difference against the number of iterations.
TMOBO consistently outperforms the alternatives across all tasks, achieving better-converged and more diversified Pareto-optimal fronts. Its trajectory-based acquisition function enables it to extract valuable insights from partially trained models. In contrast, $q$NEHVI-T and $q$EHVI-T fail to match TMOBO’s performance in maintaining convergence speed and solution quality; and ParEGO-T performs the worst. Overall, these results highlight TMOBO’s effectiveness in hyperparameter tuning scenarios, particularly in balancing computational efficiency and optimization performance.
E.3 Impact of Early Stopping
On the CNN benchmark, we further take this opportunity to provide readers with a clearer understanding of our algorithm’s early stopping mechanism as well as the computational complexity. Figure E.7 first shows that TMOBO and its non-early-stopping variant, TMOBO-nES, achieve comparable performance throughout the trial, from initialization to the end of 100 algorithm iteration. Notably, TMOBO requires less overall runtime (including CNN training time and algorithm overhead) than TMOBO-nES (as reflected by the “Total” boxplots) primarily because its early stopping mechanism prevents unnecessary training of the CNN model, thereby reducing the number of epochs per trajectory and conserving the wall-clock time. This efficiency is evident in the lower “Train” boxplot of TMOBO. The trade-off is that the early stopping mechanism introduces additional computational overhead, since TMOBO needs to determine in real time when to terminate training. However, when compared with the overall time required to train the CNN model, the overhead for both TMOBO and TMOBO-nES is negligible.
Figure E.7: Comparison of the computational efficiency of TMOBO and TMOBO-nES on the hyperparameter tuning task of MobileNetV2 on the CIFAR-10 dataset. [Left] Average log HV difference against algorithm iterations for each algorithm across 10 independent trials. The shaded region indicates two standard errors of the mean. [Right] Boxplots of wall-clock time for each algorithm, where “Total” indicates the overall runtime, “Train” the overall CNN training time, and “Overhead” the algorithm additional processing time.
E.4 Impact of Objective Complexity
In different ML applications, practitioners may consider optimizing hyperparameters for varying objectives. This makes it essential to evaluate TMOBO’s performance across diverse scenarios. To explore this, we modified the LCBench hyperparameter tuning tasks by replacing the training cost objective with validation cross-entropy, while retaining validation accuracy as the other objective. This modification introduces a more complex, non-linear relationship between objectives.
Figure E.8 shows the performance of TMOBO and three alternative algorithms in terms of the average log HV difference against the total number of epochs. As the trajectory characteristics become more complex, each algorithm takes more effort to converge. However, TMOBO consistently outperforms the alternatives in both convergence speed and final HV difference, except for the second task on “higgs” where TMOBO and qNEHVI-T demonstrate comparable performance. Despite the changes in objective complexity, TMOBO maintains its advantage over the alternatives by leveraging information from partially trained models and dynamically adjusting training durations. This capability allows it to efficiently navigate the more challenging optimization tasks.
Figure E.8: Average log Hypervolume difference against the number of epochs for each algorithm on five different hyperparameter tuning tasks (with first objective being validation accuracy and the second objective validation cross entropy). Each algorithm runs for 10 independent trials. The shaded region indicates two standard errors of the mean.
E.5 Impact of Noisy Levels
We then investigate the impact of stochasticity on the performance of TMOBO. To this end, for each synthetic problem generated from DTLZ, we add to each objective a Gaussian noise with standard deviations of 10%, 1%, and 0.1% of the objective range. Figure E.9 highlights TMOBO’s adaptability and performance across different noise scenarios.
At lower noise levels (0.1% and 1%), TMOBO demonstrates stable and consistent performance. The algorithm quickly converges to a low HV difference, showcasing its ability to reliably predict trajectories and identify Pareto-optimal trade-offs. As the noise level increased to 10%, the optimization becomes more challenging due to the amplified variability in objective evaluations, which TMOBO inherently inherits as part of the trade-off modeling. Despite these challenges, TMOBO continues to converge and achieves reasonably low HV differences by the end of trials.
Figure E.9: Average log Hypervolume difference against iterations for TMOBO on synthetic problems generated from DTLZ1 with noise level 0.1, 0.01, and 0.001. For each noise level, TMOBO runs for 10 independent trials. The shaded region indicates two standard errors of the mean.
E.6 Impact of Replication Strategy
@@
As a result of the analysis on noise levels, replicated observations at a query $(\bm{x},t)$ might be needed to achieve adequate accuracy, especially in the presence of significant noises. While this particular scenario is not the focus of our current study, we outline a straightforward method to extend the proposed TMOBO to accommodate the requirement for replicated observations when the computational budget is sufficient. It is essential to recognize that with each noisy observation obtained at $(\bm{x},t)$, the training process must start from scratch and will also output noisy observations from $(\bm{x},1)$ up to $(\bm{x},t-1)$. As a consequence, careful maintenance of multiple training procedures for replications is needed to ensure algorithmic efficiency. The pseudo-code for TMOBO-$P$, a modified version of TMOBO, is presented in Algorithm E.1.
Algorithm E.1 Framework of TMOBO-$P$
@@
Figure E.10 shows the performance of TMOBO-$P$ with $P=1$, $4$, $16$, and $64$ on the synthetic problems derived from ZDT1. TMOBO proposed in the main paper can be considered a special case of TMOBO-$P$ with $P=1$. This time, we add Gaussian noise to each objective with a standard deviation of $10\%$ of the objective range in order to emphasize the influence of noises. It can be observed that the performance of TMOBO-$P$ improves significantly in terms of the HV difference as more replications are allowed. This improvement is evident not only in rapid convergence during the initial stages but also in obtaining high-quality results at the end when comparing TMOBO with $P\geq 16$ to TMOBO with $P<16$. In the meantime, benefiting from a large number of replications, TMOBO-$64$ has stable performance and its standard error is relatively small.
Figure E.10: Average log Hypervolume difference against iterations for each algorithm on synthetic problems generated from ZDT1 with $10\%$ noise level. Each algorithm runs for 20 independent trials. The shaded region indicates two standard errors of the mean.